# Occupancy-aware Trajectory Planning for Autonomous Valet Parking in Uncertain Dynamic Environments

### v1

@@ III-C Update
The occupancy estimator combines current spot observations with predicted agent behaviors to estimate future occupancy, guiding ego vehicle navigation (Section IV). For fully observed spots, i.e., $p_{c}(d_{(\zeta,r)}(x_{t},s))=1$, we fully trust the current observation and propagate occupancy using only agent predictions. For partially observed or unobserved spots, i.e., $p_{c}(d_{(\zeta,r)}(x_{t},s))<1$, we apply Bayes’ rule to fuse the observation with the prior belief as
@@
where $x_{t}$ is the current state of the ego. Since only the current occupancy observation $Z_{t}^{s}$ is available, we set $Z_{k+1}^{s}=Z_{t}^{s}$ for all $k\in\{t,\ldots,t+T-1\}$. For all unobserved spots $s$, from (4), we have $p_{c}(d_{(\zeta,r)}(x_{t},s))\approx 0.5$}. Hence, using (16), the update equation reduces to $b_{k+1}^{s}\approx\bar{b}_{k+1}^{s}$, i.e., the prior belief remains unchanged.
@@ IV Strategy Planner
The Strategy Planner leverages probabilistic occupancy forecasts from the Spot Occupancy Estimator to select actions such as parking, waiting, or exploring, and generates a reference trajectory that balances efficiency and safety in dynamic parking environments.
IV-A Promising Spots to Park
At each planning step, we filter candidate parking spots inside the FoV based on the occupancy predictions, classifying spots that are promising for the ego vehicle to park.
$\text{Spot}~s\text{ is vacant in the future}~(t+T)\text{ if }$
@@ IV-A Promising Spots to Park
We typically set $P_{v}<P_{o}$ reflecting the asymmetric reasoning for spot availability. An initially vacant spot must remain vacant with high probability—otherwise, it is likely to be claimed by another agent and thus unsuitable for the ego to park. In contrast, an initially occupied spot can be considered promising even with a lower probability of becoming vacant, as this indicates the occupying agent is likely to leave, creating an opportunity for the ego to park. This occupancy-based filtering ensures that the strategy planner only considers realistic and feasible parking spots, optimizing decision-making efficiency and safety in dynamic environments.
IV-B Compute and Evaluate Paths
If multiple spots are estimated to be vacant within the FoV by (17), we evaluate candidate paths and prioritize them using a cost function that promotes parking efficiency and smooth maneuvers. At planning step $t$, let $\mathcal{S}_{t}\subset\mathcal{P}$ be the set of spots classified as vacant via (17). From the ego state $x_{t}=[X_{t},Y_{t},\theta_{t}]^{\top}$, we compute paths $\mathcal{T}=\{\tau_{g}\}_{g\in\mathcal{S}_{t}}$, where each $\tau_{g}$ leads to a goal state $g$. We use Hybrid A star to generate paths that avoid the static obstacles, though alternative methods (e.g., sampling or spline-based planners [13, 14]) may also be applied. Path computations are parallelized across all candidate spots $g$ for efficiency.
@@ IV-B1 Cost function
We evaluate which among $\{\tau_{g}\}_{g\in\mathcal{S}_{t}}$ is the optimal path for the ego vehicle to follow using a cost function. A path is defined as $\tau_{g}=\{x(k)\}_{k=t}^{t+N_{g}}$}, where $x(t)=x_{t}$}, $x(t+N_{g})=g$ and $N_{g}$ is the time required to reach the goal $g$ from the current state $x_{t}$. We use the following cost function to prioritize the paths.
@@
where $t_{w}$ denotes the waiting time which is explained later, $\mathcal{Y}(k)=\{\hat{y}^{j}(k)\}_{j=1}^{D}$} represents the configuration of dynamic agents at time $k$ and the static obstacles are given by $\{w_{i}\}_{i=1}^{W}$, where $w_{i}\in\mathbb{R}^{2}$. The per-stage cost $c_{o}(\cdot,\cdot)$ penalizes proximity to obstacles, while $c_{\textnormal{smooth}}(\cdot)$ promotes smooth trajectories, as defined below.
@@
$\quad\mathbbm{1}(\textrm{reverse}(k))+\mathbbm{1}(\textrm{change\_gear}(k)),$ (20)
where $\text{D}_{\text{C}}(\cdot,\cdot)$ is the cosine distance between two vectors [15]. Other terms in the smoothness cost $c_{\textnormal{smooth}}(\cdot)$ are
@@
$\mathbbm{1}(\textrm{reverse}(k))=\begin{cases}1,&\textrm{if}\ \textrm{D}_{\text{C}}\left(v(k),\begin{bmatrix}\cos(\theta(k))\\ \sin(\theta(k))\end{bmatrix}\right)>1,\\ 0,&\textrm{otherwise}\end{cases},$
$\mathbbm{1}(\textrm{change\_gear}(k))=\begin{cases}1,&\textrm{if}\ \textrm{D}_{\text{C}}\left(v(k),v(k+1)\right)>1,\\ 0,&\textrm{otherwise}.\end{cases}$ (21)
The obstacle avoidance cost $h(\cdot,\cdot)$ is
$h(x,y)=e^{-\alpha_{o}d_{o}(x,y)},$ (22)
$d_{o}(x,y)$ is the distance between the ego vehicle $x$ and obstacle $y$. The barrier cost $h(\cdot,\cdot)$ penalizes proximity to obstacles, with $\alpha_{o}=2$ for dynamic agents and $\alpha_{o}=3$ for static ones, reflecting smaller penalties for static obstacles. We extend the $3$-circle distance from [16] to a $5$-circle model for $d_{o}(\cdot,\cdot)$ to improve accuracy for dynamic collision avoidance. For static obstacles, we instead compute the exact distance between the ego vehicle’s edge and obstacle points $\{w_{i}\}_{i=1}^{W}$ [9], since static obstacle positions are more reliably perceived. To balance terms, $\|a(t)\|$ and the cosine distance are scaled to $[0,1]$ using velocity limits, while $h(x,y)\in[0,1]$ from (22).
@@ IV-B2 Waiting time
We incorporate wait and go behaviors into the planned paths $\tau_{g}$ using the waiting time $t_{w}$ in (18). This enables the ego vehicle to yield to dynamic agents and park in a spot $g$ that is estimated to be vacant in the near future by our Spot Occupancy Estimator (Section III). To ensure feasibility, we impose a maximum wait time $T_{w}$ and evaluate only paths with $t_{w}\leq T_{w}$.
Remark: If $T<N_{g}$, we extrapolate the predictions using a constant velocity model as $\hat{y}^{j}(k)=\int_{s=t+T}^{k}\hat{v}^{j}(t+T-1)ds$} for all $k\geq t+T+1$} and $j\in\{1,2,\ldots,D\}$}, where $\hat{v}^{j}(k)=\frac{\hat{y}^{j}(k+1)-\hat{y}^{j}(k)}{\Delta t}$}.
Once we compute the paths $\mathcal{T}=\{\tau_{g}\}_{g\in\mathcal{S}_{t}}$ and their associated costs $\mathcal{C}(\tau_{g})$, we choose the path that has the minimum cost $\operatorname*{arg\,min}_{\tau\in\mathcal{T}}\mathcal{C}(\tau)$ and avoids all the obstacles respecting the safety thresholds. We set larger safety thresholds for dynamic agents—$50$ cm for vehicles and $90$ cm for pedestrians—compared to $20$ cm for static obstacles.
@@ IV-C Exploration
If no feasible path exists to a spot in $\mathcal{S}_{t}$, the ego vehicle switches to exploratory mode, selecting a path that maximizes information gain about spot occupancy. In typical parking lots such as Fig.1, maneuver options are limited: continuing straight within a row, entering a row from outside, and turning left or right at its end. We generalize this decision-making process by evaluating paths to a finite set of exploration goals $\mathcal{E}\subset\mathbb{R}^{3}$, where each $e\in\mathcal{E}$ denotes a candidate goal pose.
@@ IV-C1 Exploration goals
We define three exploration goals $\mathcal{E}=\{e_{1},e_{2},e_{3}\}$} at each time step: going straight ($e_{1}$), turning left ($e_{2}$), and turning right ($e_{3}$). These options suit structured parking lots with parallel rows (Fig. 1), though we plan to extend them to more general layouts such as multi-level garages in future work. The straight goal is
@@
where $x_{t}=[X_{t},Y_{t},\theta_{t}]^{\top}$ is the ego state, $\epsilon$ is the full observability threshold (Sec. II-C), and $0\leq\eta<\epsilon$ denotes how far the ego moves forward.
Left and right turn goals are
$e_{2}$ $=\begin{bmatrix}R(\theta_{t})\begin{bmatrix}x_{\textnormal{road}}+\frac{l_{w}}{2}\\ \epsilon\end{bmatrix}+\begin{bmatrix}X_{t}\\ Y_{t}\end{bmatrix}\\ \\ \theta_{t}+\frac{\pi}{2}\end{bmatrix}$, (24)
$e_{3}$ $=\begin{bmatrix}R(\theta_{t})\begin{bmatrix}x_{\textnormal{road}}-\frac{l_{w}}{2}\\ -\epsilon\end{bmatrix}+\begin{bmatrix}X_{t}\\ Y_{t}\end{bmatrix}\\ \\ \theta_{t}-\frac{\pi}{2}\end{bmatrix}$,
where $x_{\textnormal{road}}$ is the center of the newly observed road, and $l_{w}$ its lane width. The offsets $\pm\frac{l_{w}}{2}$ encode left and right turns under right-hand traffic convention, with corresponding adjustments to heading and $y$-coordinate sign.
@@ IV-C2 Information gain
We generate paths to the exploration goals using Hybrid A star for $e_{2}$ and $e_{3}$, and a 5th-order spline to interpolate a path from the current state $x_{t}$ to $e_{1}$. Let $\mathcal{T}^{\prime}=\{\tau_{e}\}_{e\in\mathcal{E}}$ denote the set of paths from $x_{t}$ to each goal. To evaluate the information gain of a path, we consider the entropy of a parking spot $s$ at time $k$ as
@@
where $b_{k}^{s}=P(O_{k}^{s}=1|Z_{k}^{s})$ is the probability that spot $s$ is occupied given observation $Z_{k}^{s}$. We simulate future observations $\{Z_{k}^{s}\}_{k=t+1}^{t+T}$ along each path $\tau_{e}$ and denote $Z_{k}^{s}(\tau_{e})$ to be the observation of spot $s$ at time $k$ when the ego follows path $\tau_{e}$. Then, we recursively update beliefs $b_{k}^{s}(\tau_{e})$ using the update rule (16) and the new simulated observations $Z_{k}^{s}(\tau_{e})$ for each path $\tau_{e}$. The total information gain of path $\tau_{e}$ for all parking spots is defined to be the entropy reduction from following the path:
@@
where $N_{e}$ is the length of path $\tau_{e}=\{x(k)\}_{k=t}^{t+N_{e}}$. To balance exploration with motion efficiency, we use (18) and (26) to define the exploration cost
@@
and select the optimal path as $\operatorname*{arg\,min}_{\tau_{e}\in\mathcal{T}^{\prime}}\mathcal{C}_{e}(\tau_{e})$, subject to safety thresholds. If no safe path exists, a backup planner keeps the vehicle stationary or steers it away from adversarial dynamic agents on a collision course.
@@ V Experimental Results
The task for the ego vehicle is to navigate from the initial position to a feasible parking spot, assuming that the onboard sensors provide real-time information on static vehicles, driving aisles and dynamic agents within the FoV. We evaluate our framework in a simulated parking lot environment designed to emulate a shopping mall layout as shown in Fig. 1. The mall entrance is located at the top edge, centered along the $x$-axis. To reflect realistic congestion patterns, the initial occupancy probability of each parking spot is randomized using a binomial distribution, with higher occupancy near the entrance and lower occupancy farther away as presented in Fig. 3. For evaluation, we conduct ablation studies and baseline comparisons with existing valet parking approaches, and compute key performance metrics in Table III. We run 50 randomized experiments for each method in Table III, randomizing the parking lot configuration, ego vehicle’s initial position, and the behavior of other dynamic agents. The ego vehicle is initialized either inside a parking row or outside, while other vehicles may start inside vacant spots or outside. The paths of other vehicles are generated using a local Hybrid A star planner for each vehicle entering or exiting a spot. Dynamic pedestrians are spawned randomly outside parking spots and move with velocities $v\sim\mathcal{U}[-1.5,1.5]$ m/s. All simulations are performed in Python 3.13 on Ubuntu 20.04 with Intel Xeon E5-2643 v4 CPU.
Simulation parameters: The vehicle dynamics are modeled using the kinematic bicycle model (1), discretized with a time step of $0.1$ s for simulation. The vehicle parameters are listed in Table I, based on the 2024 Honda Accord [17]. Our Hybrid A planner employs a state grid resolution of $2$ m in both $X$ and $Y$ directions, and $10\ \textrm{deg}$ for heading. The steering input is discretized into 5 values over $\delta_{\textrm{max}},\delta_{\textrm{max}}]$, while velocity is discretized into 5 values over $v_{\textrm{max}},v_{\textrm{max}}]$. Parking lot parameters are adopted from [18] and given in Table II. We use a prediction horizon of $T=5$ s for our spot occupancy estimator and a maximum wait time of $T_{w}=5$ s in our strategy planner. The arrival rate $\lambda_{a}$ is 0.000624 (14), while the departure rate $\lambda_{d}$ is 0.000378 (12), both of which are obtained from the parking lot data in [19]. The probability thresholds in (17) are $P_{v}=0.3$ and $P_{o}=0.7$. The scaling factor is $r=\begin{bmatrix}(2.5V_{L})^{-1}&(3V_{W})^{-1}\end{bmatrix}^{\top}$ and forward shift is $\zeta=0.5V_{L}$ in (2), while $\eta=0.5$ in (23). The observability thresholds are $\epsilon=1$ and $\gamma=1.5$.

### v2

@@ III-C Update
The update combines current spot observations with $\bar{b}_{k+1}^{s}$ to estimate future occupancy. For fully observed spots, i.e., $p_{c}(d_{(\zeta,r)}(x_{t},s))=1$, we completely trust the current observation and propagate occupancy using only agent predictions. For partially observed or unobserved spots, i.e., $p_{c}(d_{(\zeta,r)}(x_{t},s))<1$, we apply Bayes’ rule to fuse the observation with the prior belief:
@@
where $x_{t}$ is the current ego state. Since only the current occupancy observation $Z_{t}^{s}$ is available, we set $Z_{k+1}^{s}=Z_{t}^{s}$ for all $k\in\{t,\ldots,t+T-1\}$. For unobserved spots $s$, from (4), we have $p_{c}(d_{(\zeta,r)}(x_{t},s))\approx 0.5$}. Hence, using (16), the update equation reduces to $b_{k+1}^{s}\approx\bar{b}_{k+1}^{s}$, i.e., the prior belief remains unchanged.
@@ IV Strategy Planner
Given the probabilistic forecasts from the spot occupancy estimator (Sec. III), the strategy planner selects actions (park, wait, or explore) and generates a reference trajectory that balances efficiency and safety in dynamic parking scenarios.
IV-A Filter Spots
At each planning step, we filter candidate spots inside the FoV based on the occupancy predictions, considering two cases: (i) initially vacant and remain vacant (ii) initially occupied and become vacant. Formally, spot $s$ is vacant in the future (t+T) if:
@@ IV-A Filter Spots
This filtering ensures that the strategy planner only considers feasible parking spots, enhancing decision-making efficiency in dynamic environments. We set a spot is more likely to be occupied if initially vacant ($P_{v}<P_{o}$) and vice-versa to reflect the asymmetry in availability reasoning.
IV-B Generate and Evaluate Paths
At planning step $t$, let $\mathcal{S}_{t}\subset\mathcal{P}$ denote spots classified as vacant by (17). We generate trajectories $\mathcal{T}=\{\tau_{g}\}_{g\in\mathcal{S}_{t}}$} using Hybrid A [^1] Alternatives include sampling or spline-based planners [32, 33, 8]., where each $\tau_{g}$ leads to a goal state $g$ from the ego state $x_{t}$. Each trajectory is defined as $\tau_{g}=\{x(k)\}_{k=t}^{t+N_{g}}$}, where $x(t)=x_{t}$}, $x(t+N_{g})=g$ and $N_{g}$ is the time required to reach $g$ from $x_{t}$. While static obstacles are considered in trajectory generation, dynamic obstacles are handled via the cost function and waiting time parameter, defined in the next section. Trajectory generations and evaluations are parallelized across all goal states $g$ for efficiency.
@@ IV-B1 Cost function
The cost function primarily evaluates efficiency (in time and distance), safety (w.r.t. collision avoidance), and smoothness (w.r.t. acceleration and gear changes):
@@
where
@@
$\quad\mathds{1}(\textrm{reverse}(k))+\mathds{1}(\textrm{change\_gear}(k)),$ (20)
$t_{w}$ denotes the waiting time (explained in the next section), $\mathcal{Y}(k)=\{\hat{y}^{j}(k)\}_{j=1}^{D}$} represents the configuration of dynamic agents at time $k$ and the static obstacles are given by $\{w_{i}\}_{i=1}^{W}$ with $w_{i}\in\mathbb{R}^{2}$. At each time step, $c_{o}(\cdot,\cdot)$ penalizes proximity to obstacles and $c_{\textnormal{smooth}}(\cdot)$ promotes smooth trajectories.
In $c_{o}(\cdot,\cdot)$, the obstacle avoidance cost $h(\cdot,\cdot)$ is
$h(x,y)=e^{-\alpha_{o}d_{o}(x,y)},$ (21)
where $d_{o}(x,y)$ is the distance between the ego vehicle $x$ and obstacle $y$. The barrier cost $h(\cdot,\cdot)$ penalizes proximity to obstacles, with $\alpha_{o}=2$ for dynamic agents and $\alpha_{o}=3$ for static obstacles, reflecting lower penalties for static ones. We extend the 3-circle distance from [34] to a 5-circle model for $d_{o}(\cdot,\cdot)$ to improve accuracy in dynamic collision avoidance. For static obstacles, the exact distance between the ego vehicle’s edge and obstacle points $\{w_{i}\}_{i=1}^{W}$ is computed as given in Section III of [26], since static obstacles are more reliably perceived. To balance terms, $\|a(t)\|$ and the cosine distance are scaled to $[0,1]$ using velocity limits.
In (22), $\text{D}_{\text{C}}(\cdot,\cdot)$ is the cosine distance [35] and
@@
$\mathds{1}(\textrm{reverse}(k))=\begin{cases}1,&\textrm{if}\ \textrm{D}_{\text{C}}\left(v(k),\begin{bmatrix}\cos(\theta(k))\\ \sin(\theta(k))\end{bmatrix}\right)>1,\\ 0,&\textrm{otherwise}\end{cases},$
$\mathds{1}(\textrm{change\_gear}(k))=\begin{cases}1,&\textrm{if}\ \textrm{D}_{\text{C}}\left(v(k),v(k+1)\right)>1,\\ 0,&\textrm{otherwise}.\end{cases}$ (22)
@@ IV-B2 Waiting time
We incorporate wait-and-go behaviors in the trajectories $\tau_{g}$ by introducing a waiting time $t_{w}$ in (18). This enables the ego vehicle to pause and yield to dynamic agents before moving to park in a spot $g$, thereby reducing collision risks in constrained parking lots. To ensure feasibility, we impose a maximum wait time $T_{w}$ and consider only paths with $t_{w}\leq T_{w}$. The value of $T_{w}$ is set based on parking lot density, with future work aimed at dynamically adapting it to congestion.
Remark: If $T<N_{g}$, we extrapolate the predictions using a constant velocity model: $\hat{y}^{j}(k)=\int_{s=t+T}^{k}\hat{v}^{j}(t+T-1)ds$} for all $k\geq t+T+1$} and $j\in\{1,2,\ldots,D\}$}, where $\hat{v}^{j}(k)=\frac{\hat{y}^{j}(k+1)-\hat{y}^{j}(k)}{\Delta t}$}.
Once we compute and evaluate the trajectories $\mathcal{T}$, we choose the one that has the minimum cost $\operatorname*{arg\,min}_{\tau\in\mathcal{T}}\mathcal{C}(\tau)$ and avoids all obstacles respecting the safety thresholds. We set larger safety thresholds for dynamic agents—5050cm for vehicles and 9090cm for pedestrians—compared to 2020cm for static obstacles. These values are tailored to parking scenarios but can be adjusted for other applications.
@@ IV-C Exploration
If no feasible trajectory exists to a spot in $\mathcal{S}_{t}$, the ego vehicle switches to exploratory mode, selecting a trajectory that maximizes information gain about spot occupancy. In typical parking lots like Fig. 2, maneuver options are limited to continuing straight within a row, entering a row from outside, or turning left or right at the end. We generalize this by evaluating paths to a finite set of exploration goals $\mathcal{E}\subset\mathbb{R}^{3}$, where each $e\in\mathcal{E}$ represents a candidate goal pose.
@@ IV-C1 Exploration goals
We define three exploration goals $\mathcal{E}=\{e_{1},e_{2},e_{3}\}$} at each time step: going straight ($e_{1}$), turning left ($e_{2}$), and turning right ($e_{3}$). These options suit structured parking lots with parallel rows (Fig. 2), though we plan to extend them to more general layouts such as multi-level garages in future work. The straight goal is:
@@
where $x_{t}=[X_{t},Y_{t},\theta_{t}]^{\top}$ is the ego state, $\epsilon$ is the full observability threshold (Sec. II-C), and $\eta\in[0,\epsilon)$ denotes how far the ego moves forward. Left and right turn goal states for right-hand traffic convention are described as
$e_{\sigma}=\begin{bmatrix}R(\theta_{t})\begin{bmatrix}x_{\text{road}}+\sigma\cdot\left(\frac{l_{w}}{2}\right)\\ \sigma\cdot\epsilon\end{bmatrix}+\begin{bmatrix}X_{t}\\ Y_{t}\end{bmatrix}\\ \theta_{t}+\sigma\cdot\left(\frac{\pi}{2}\right)\end{bmatrix},\quad\sigma\in\{+1,-1\},$ (24)
where $e_{2}=e_{+1}$(left), $e_{3}=e_{-1}$(right), $x_{\textnormal{road}}$ is the center of the newly observed road, and $l_{w}$ its lane width.
@@ IV-C2 Information gain
We use Hybrid A to generate trajectories to $e_{2}$ and $e_{3}$, and a 5th-order spline to interpolate from $x_{t}$ to $e_{1}$. Let $\mathcal{T}^{\prime}=\{\tau_{e}\}_{e\in\mathcal{E}}$ denote the set of trajectories to each exploration goal. The information gain of a trajectory is evaluated using the entropy $\mathcal{H}(\cdot)$ of a spot $s$ at time $k$:
@@
where $b_{k}^{s}=P(O_{k}^{s}=1|Z_{k}^{s})$. We simulate future observations $\{Z_{k}^{s}\}_{k=t+1}^{t+T}$ for each trajectory $\tau_{e}$ and denote $Z_{k}^{s}(\tau_{e})$ to be the observation of spot $s$ at time $k$ when the ego follows $\tau_{e}$. Then, we recursively update beliefs $b_{k}^{s}(\tau_{e})$ using the update rule (16) and the simulated observations $Z_{k}^{s}(\tau_{e})$ for each $\tau_{e}$. The total information gain of a trajectory is defined using the entropy reduction of all parking spots when following $\tau_{e}$:
@@
where $N_{e}$ is the length of $\tau_{e}=\{x(k)\}_{k=t}^{t+N_{e}}$. To balance exploration with motion efficiency, we use (18) and (26) to define the exploration cost
@@
and select the optimal trajectory as $\operatorname*{arg\,min}_{\tau_{e}\in\mathcal{T}^{\prime}}\mathcal{C}_{e}(\tau_{e})$, subject to safety thresholds. If no safe trajectory exists, a contingency planner keeps the vehicle stationary or steers it away from adversarial dynamic agents on a collision course.
@@ V Experimental Results
The ego vehicle’s task is to navigate from its initial position to a feasible parking spot, assuming real-time onboard sensor data on static vehicles, driving aisles, and dynamic agents within the FoV. We evaluate our framework in a simulated shopping mall parking lot (Fig. 2), with randomized initial spot occupancy via a binomial distribution reflecting higher congestion near the entrance (Fig. 4). We conduct ablation studies and comparisons with existing valet parking methods that are summarized in Table III, with 50 randomized runs per method varying initial occupancy, ego start position, and dynamic agent behavior. The ego is initialized either inside a parking row or outside, while other vehicles may start in vacant spots or outside, with their paths generated by a local Hybrid A planner for exiting or entering spots, respectively. Dynamic pedestrians are randomly spawned outside the occupied spots with velocities $v\sim\mathcal{U}[-1.5,1.5]$ m/s. All simulations are run in Python 3.13 on Ubuntu 20.04 with an Intel Xeon E5-2643 v4 CPU.
Simulation parameters: The vehicle dynamics are modeled using the kinematic bicycle model [27], discretized with a time step of $0.1$ s for simulation. The vehicle parameters are listed in Table I based on the 2024 Honda Accord [36]. Our Hybrid A planner employs a state grid resolution of $0.2$ m in both $X$ and $Y$ directions, and $10\ \textrm{deg}$ for heading. Steering and velocity inputs are each discretized into 5 values over their limits $\delta_{\textrm{max}},\delta_{\textrm{max}}]$ and $v_{\textrm{max}},v_{\textrm{max}}]$, respectively. Parking lot parameters in Table II are adopted from [37]. The prediction horizon is $T=5$ s with a discretization step of $0.1$ s, and the maximum wait time is $T_{w}=5$ s. The average arrival and departure rates are  $\lambda_{a}=0.000624~\textrm{s}^{-1}$ (14) and $\lambda_{d}=0.000378~\textrm{s}^{-1}$ (12), respectively, obtained from [30]. The probability thresholds in (17) are $P_{v}=0.3$ and $P_{o}=0.7$. The scaling factor is $r=\begin{bmatrix}(2.5V_{L})^{-1}&(3V_{W})^{-1}\end{bmatrix}^{\top}$ and forward shift is $\zeta=0.5V_{L}$ in (2), while $\eta=0.5~\textrm{m}$ in (23). The observability thresholds are $\epsilon=1$ and $\gamma=1.5$, and $\alpha_{c}=25$ in (4).